SX LUCI/AO Commissioning

01:20 Telescope Open

01:45 Installing new Collimation model

02:56 Closed loop bin2
  • opt gain = 2.55, 2.0. 0.82
  • too high. set to 2.0, 1.25, 0.82
  • TN 025805

  • TN 030441 movie bump_at_2

  • Edge at 2:00 rolled when actuator 618 jumped. 618 is 5th row in at 2:24

03:15 Put 618 in pos do-not-use list. Put 560(? from the 22nd) back in use.
  • Large current on 618 even though it is in pos-do-not-use list
  • Only 220 mode
  • IMAGE
  • put in current do-not-use list

03:35 Flat with 655 modes

03:44 Close loop in bin 1
  • no got
  • 034411

03:48 Closed loop bin 1, 400 mode
  • Opt gain, Get rolled edge at 2 to 3
    • movie bin_1_opt_gain_bump_2.3
  • 2nd try Opt gain, Get rolled edge at 1 and 5
    • movie bin_1_opt_gain_bump_1.5

  • 041818 10000 frames
  • skip fram

  • Seeing 0.9 on DIM
  • 042845 OK
  • 049901 Rolled edge at 1:00

  • 0437829 High currents in center of shell at 8:00
  • Safe skip and then open

04:45 bin 2 preset
  • closed loop
  • Optimize gain. skip frame during HO2. caught in TN 045046
  • Seeing 1.28
  • Loop closed with LNgains.
  • Get skip frames and loop open

05:04
  • bin 2 closed loop with LN gains
  • Skip frame caught in TN 050422

07:00 Change auto gain.conf
  • copy original to auto_gain.conf.20170924
  • Change bin*_max_tt, _ho1, _ho2 to lower value

07:05 Preset to bin 2 star
  • optimize gain worked
  • gains 1.7, 1.5, 0.97
  • Loop closed for 10 minutes. No safe skips
  • Well, cloud came over and photons went to 0. Loop opened

07:25 Preset to bin 1 star
  • high clouds forming overhead. Loose light. AcquireRefAO failed
  • better so send runAO again
  • success.
  • optimize gain 1.27, 0.88, 0.2
  • 0741??
  • gain 1.27, 0.88, .0.45
  • Gopt on, NCPA on
  • 074532

08:17
  • Dark 081739
  • 081918 no luci
  • 082100, 082141, 082206with LUCI FeII (almost situated)
  • 082141
  • 082206
  • Move away from bad pixel
  • 082414, 082519, 082536, 082638, 082811

09:38 Jochen send LUCI Preset
  • bin 2, 10.57 mag
  • Success
    • Gopt = 0.1
  • Luci take data

08:51 Rolled an edge at 11:30. Loop open
  • closed loop manually
  • opt gain = 1.7, 1.5, 1.0
  • 085556

09:13 Safe skip loop opens
  • Act 217 on center ring has high force so only 384 modes

10:15 Preset to bin 1 star (ends up bin 2 with 10.5 mag)

11:27 act 196 failed

12:20 ESM with bin 4
  • reconstructors or bin 2 and 4 are not correct in table_LUCIFER_TTM-AO.txt

-- %USERSIG{DougMiller - 2017-09-24}%

Topic revision: r1 - 25 Sep 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback